On the basis of the following data, determine the value of the inventory at the lower of cost or market. Assemble the data in the form illustrated in Exhibit 9.

Inventory
Item

Inventory
Quantity

Cost per
Unit

Market Value per Unit
(Net Realizable Value)

Birch 44 $53 $31
Cypress 33 51 55
Mountain Ash 33 152 149
Spruce 33 277 275
Willow 41 295 291
Inventory at the Lower of Cost or Market
Inventory Item Total Cost Total Market Total Lower of C or M
Birch $ $ $
Cypress
Mountain Ash
Spruce
Willow
Total $ $ $
Answer #1
Inventory at the Lower of Cost or Market
Inventory Item Total Cost Total Market Total Lower of C or M
Birch 2332 1364 1364
Cypress 1683 1815 1683
Mountain Ash 5016 4917 4917
Spruce 9141 9075 9075
Willow 12095 11931 11931
Total 30267 29102 28970
Formulas:
Inventory Item Total Cost Total Market
Birch =44*53 =44*31
Cypress =33*51 =33*55
Mountain Ash =33*152 =33*149
Spruce =33*277 =33*275
Willow =41*295 =41*291